Passing through (-2,-7) and parallel to the line whose equation is y = -5x + 4
can you please explain to me a little better my instructor is all over the place.

Answer by stanbon(75887)

Passing through (-2,-7) and parallel to the line whose equation is y = -5x + 4
-------------
The given line has a slope = -5
Any line parallel to it must have a slope of -5
-----------------------------
Equation form: y = mx+b
--------
m = -5 and y = -7 when x = -2.
Solve for "b":
-----
-7 = -5*-2 + b
-7 = 10 + b
b = -17
-------------
Equation you want:
y = -5x - 17
=========================
Cheers,
Stan H.